### v12

@@ Abstract
Data from national accounts support an argument that capital growth and acceleration arrive without help from net saving. We explore ways in which this is possible, and discuss implications for economic teaching and public policy.
@@ 1 Introduction and overview
Many economists over the centuries have reasoned that net saving [^1] As reported in national accounts; saving and investment differ only by statistical discrepancy. We view them as different, for reasons shown in Appendix A, but will equate them in this paper by using the word ”saving” in the Keynesian sense of invested saving. should tend to give equal capital growth. Economists since the early nineteenth century have added the proviso that net saving cannot safely outpace innovation; more capital must mean capital redesigned for greater productivity if economies are to escape risk of capital glut and diminishing returns (West (1815), Ricardo (1815), Malthus (1815)). Roy Harrod (1939)) described that limit for safe net saving, meaning the rate of imagining and developing new ideas for more productive forms of capital, as the “warranted rate”. Harrod, and many other economists of his time and since, have focused on growth of output rather than of capital, but have modeled growth of output by first assuming the equivalence of net saving and capital growth, within the warranted rate, and then looking for effects of that capital growth on later output growth.
@@
We test between those two explanations of capital growth, by net saving or by increase in productivity of capital and labor already in existence, by comparing net saving to concurrent change in market-value capital in 92 countries, and also comparing changes in those two variables. All data are drawn from national accounts of those countries as collated on the free website World Inequality Database. Our findings are easiest to reconcile in an interpretation that net saving adds nothing to capital growth. [^2] We argue, however, that productivity gain will not give equal capital growth unless gross saving is enough to offset depreciation. We interpret any shortfall in this coverage as a negative component in net saving, and expect capital growth to equal productivity gain less that deficit (see Section 3). Net saving, if so, raises the physical quantity of capital, but not the aggregate value, and so reduces the value per unit. This condition describes capital glut.
The data are most easily explained by the present value principle, and by production efficiencies enabled through innovation. Value is created in the mind of the market at the moment when prospective cash flows are discounted. It is created only if the market sees a path, step by step, from the start, to practical realization of those prospective cash flows. Then capital growth arrives when the market first evaluates prospective cash flows, and is realized eventually in physical outcomes insofar as the market has predicted correctly. That is to say that capital growth, in the mind of the market, precedes and pays for any needed investment in new plant and equipment. [^3] Present value is time-discounted, and affords innovators only enough, at first, for first steps in implementation. As implementation proceeds, present value rises with time and affords more. Early investors borrow from the future, in effect, and pay the interest from increments in present value as positive cash flows approach with time.
@@
Meanwhile the innovator acquires materials and plant capacity and labor skills at market prices determined by their uses in current technology, but applies them more productively until competition catches up. It is that temporary market advantage to the innovator, which we call the innovator’s reserve, that explains capital growth without net saving in a practical and mechanical sense, while the present value principle gives the source of investment. This idea will be called free growth theory for easy reference.
@@
It predicts only at the largest scales, and only for the private sector. Individuals and groups and even small economies can grow in capital through saving if others dissave equally. That possibility is foreclosed only at the scale of all capital and all economies together. The public sector, meanwhile, responds to political rather than market choices, and grows or shrinks accordingly.
@@
Inferences for economic teaching include the obvious ones for growth theory and for net saving in general. They include others as well. One of the central doctrines of the marginalist revolution has held that market realization converges to producer cost, when that cost includes imputed interest on assets owned. Net saving gives producer cost, and falls short of market realization in the presence of technological growth from new ideas. Meanwhile the doctrine that net income equals consumption plus net saving is put into question by evidence offered here suggesting that net saving increases the physical quantity of capital, but not the aggregate value.
@@ 2 Net saving, capital growth and capital acceleration
Thrift theory will mean the combined ideas that net saving $S_{net}$ is realized in equal growth of capital $K$, if $S_{net}$ holds within the warranted rate, and that there is no other source of capital growth except market noise regressing to zero. "Thrift assumptions" will mean those two plus the assumptions that current $S_{net}$ holds within the warranted rate. The acronym uta will mean "under thrift assumptions." Then
@@
Test results for the predictions $\frac{s^{*}}{g(K)}=\theta=1$, under thrift assumptions, appear in Fig. 1 and in Tables 1, 2 and 3. Screens to which they refer are explained in Section 16 below. Results were found as GDP-weighted averages over all countries and years, from data for net saving and market-value capital taken from national accounts. Figures 1 and 2 show $\frac{s^{*}}{g(K)}$ and $\frac{\Delta s^{*}}{\Delta g(K)}$ as 0.481 and 0.064 respectively. Table 1 shows regression of $s^{*}$ on $g(K)$ as 0.0771, and regression of $\Delta s^{*}$ on $\Delta g(K)$ as 0.0559.
Table 1: Regression of $s^{*}$ on $g(K)$ and $\Delta s^{*}$ on $\Delta g(K)$, GDP-weighted. Screen = 0.01. $H_{0}$ per thrift theory: $\frac{s^{*}}{g(K)}=\frac{\Delta s^{*}}{\Delta g(K)}=1$.
@@
Table 2: Average $\frac{s^{*}}{g(K)}$ in 92 countries (screen = 0.01). Number of years clearing screen shown in ()
@@
Note: $s^{*}$ is defined as $\frac{S_{net}}{K}$. Thrift theory predicts $\frac{s^{*}}{g(K)}=1$. Free growth theory makes no prediction for these data.
Table 3: Average $\theta$ in 92 countries (screen = 0.01). Number of years clearing screen shown in ()
@@
Note: $\theta$ is defined as $\frac{\Delta s^{*}}{\Delta g(K)}$. Thrift theory predicts $\theta=1$. See Section 3 for predictions of free growth theory.
3 Interpretation of test results
The finding $\frac{s^{*}}{g(K)}=0.481$ reveals average observed proportions between $s^{*}$ and $g(K)$, we repeat, and does not of itself show causal relatedness between those two variables. Causal relatedness is shown rather by the degree of constancy in proportions, and thus on regression of $s^{*}$ on $g(K)$ found at 0.0771.
$\theta$ or $\frac{\Delta s^{*}}{\Delta g(K)}$, measured at $0.064$, gives a separate measure of causal relatedness between $s^{*}$ and $g(K)$ as explained. Regression of $\Delta s^{*}$ on $\Delta g(K)$, found at 0.0559, gives a measure of causal relatedness between $\Delta s^{*}$ and $\Delta g(K)$. Might these small but positive findings allow the possibility that some capital growth is explained by net saving?
Free growth theory predicts that the present value of new ideas can pay for equal capital growth, but only after gross saving has made up for all depreciation. In years of recession, households may defer part or all of depreciation saving to support consumption. In these years, growth in the present value of better ideas will not realize equal net growth of capital, as some of it must offset any shortfall in depreciation saving first. Allowing for these years, we can say more compactly that free growth theory predicts that a gain in the present value of new ideas accounts for all capital creation, and thus for the sum of capital growth and restoration of shortfalls in depreciation saving. Thus free growth theory does not predict a zero value for $\theta$, but rather predicts $\theta$ to equal the share of capital creation applied to restoring those shortfalls. With this understanding, free growth theory could fit the appearance in Fig. 2 that $\theta$ found higher values during the period of the Iraq-Kuwait, dot com and subprime crises. We will suppose that 0.064 gives a plausible estimate of the shortfall in depreciation saving, until more is known, and will proceed on the simplifying assumption that free growth theory predicts correctly.
4 Inferences for the creative destruction idea
Creative destruction (Schumpeter (1912) [^4] Published by Schumpeter in 1912, and Translated by Opie in 1934.) is obsolescence of existing capital through innovation of new and preferred technology, as with horse-drawn carriages and automobiles. Granting the obsolescence of old ways by new, it does not follow that innovation reduces values of assets already present. Data cited in this paper appear to show that capital growth is maximized when investment is enough to offset depreciation. A natural source of that investment is depreciation of assets the new technology will replace, and a natural market for the new technology may be owners of the assets to be replaced. Carriage owners were better able to afford cars when the carriages in which they had invested before lived out their economic lives as originally expected, so that the depreciation component in the cash flows originally expected can pay for cars in replacement just as it might have paid for new carriages.
This analysis suggests that market forces favor a kind through orderly obsolescence through replacement without premature displacement as the new technology awaits its turn.
@@ 5 Solow’s puzzle
Solow was right in that much innovation requires months or years of gross investment to pay for negative cash flows in building new plant before the new products roll out from the factory doors. What he left unsaid is that this investment would not have been forthcoming unless positive cash flows at the end of the tunnel were predicted to be sufficient to create positive present value from the start. Data in our displays suggest that this initial creation of value, plus depreciation investment, are enough.
@@ 6 Mechanics of free growth
The present value argument just given offers an accounting explanation of embodied growth in terms of sources of funds. The step-by-step mechanical explanation, we suggest, in terms of things rather than funds, is that embodied growth is disembodied growth on a finer scale. At each step toward realization of the new plant and products, raw materials and products and labor skills and plant capacity currently available on the market are adapted to new uses. The innovator pays for these inputs at a market price determined by their value in established productive uses, but applies them innovatively to realize higher prospective cash flows, and hence higher present values, to the innovator (Marshall (1890), Schumpeter and Opie (1934)). This difference in present value realized less price paid will here be called the “innovator’s reserve”, meaning reserve price for inputs of capital and labor.666i.e., capital and labor inputs are worth more to the innovator in that the innovator applies them in ways to realize greater returns. The present value of additional cash flow enabled by this advantage in return quantifies the innovator’s reserve, and equivalently the non-random component of free growth. The innovator’s reserve quantifies the part of free growth explained by productivity gain as distinct from random market noise. As such, it is the quantity added to depreciation saving to enable embodied growth, so that net saving is never needed.
Our findings support those of Piketty and Zucman (2014) and Kurz (2023) as to the market power of innovators to explain capital growth beyond net saving. Those studies assumed that net saving is recovered in capital growth, so that this market power, which we call the innovator’s reserve, need explain only the part of capital growth remaining. Again, we go farther by questioning the assumption that net saving contributes even a part of capital growth.
@@ 7 Tech Externalities
The innovator’s reserve, quantified by the present value of the innovator’s idea, gives the amount of free growth to the innovator. Others may benefit indirectly. Old factories and warehouses and residences gain in utility as they house higher-tech equipment and worker skills. Old highways do the same as they carry higher-tech freight and drivers and passengers to higher-tech destinations in higher-tech vehicles. [^7] Rae (1834) makes essentially the same argument in terms of the technology of his time.
Most externalities, [^8] Externalities are defined as effects of events on nonparticipants. positive and negative, are part of random market noise expected to offset to zero over time and scale. The examples above, which we call tech externalities, can impose a prevailingly positive vector resisting the downward pull of depreciation. Free growth, in the full sense, means effects of the innovator’s reserve on capital of the innovator, plus effects of tech externalities on capital in general.

### v13

@@ Abstract
Data from national accounts suggest that net saving, at the collective scale, contributes less than 8% to growth in capital value. We explore ways in which this is possible, and discuss implications for economic teaching and public policy.
@@ 1 Introduction and overview
Many economists over the centuries have reasoned that net saving [^1] As reported in national accounts; saving and investment differ only by statistical discrepancy. We view them as different, for reasons shown in Appendix A, but will equate them in this paper by using the word ”saving” in the Keynesian sense of invested saving. should tend to give equal capital growth. Economists since the early nineteenth century have added the proviso that net saving cannot safely outpace innovation; more capital must mean capital redesigned for greater productivity if economies are to escape risk of capital glut and diminishing returns (West (1815), Ricardo (1815), Malthus (1815)). Roy Harrod (1939) described that limit for safe net saving, meaning the rate of imagining and developing new ideas for more productive forms of capital, as the “warranted rate”. Harrod, and many other economists of his time and since, have focused on growth of output rather than of capital, but have modeled growth of output by first assuming the equivalence of net saving and capital growth, within the warranted rate, and then looking for effects of that capital growth on later output growth.
@@
We test between those two explanations of capital growth, by net saving or by increase in productivity of capital and labor already in existence, by comparing net saving to concurrent change in market-value capital in 92 countries, and also comparing changes in those two variables. All data are drawn from national accounts of those countries as collated on the free website World Inequality Database. Our findings imply that the efficiency of net saving in realizing capital growth has been less than 8%. We will argue that it may be substantially less, and possibly zero. [^2] We will find three measures, of which the highest is 7.71%, of the causal relatedness between net saving and capital growth. These measures do not give the direction of causality. We will argue that exogenous disruptions such as wars or epidemics or supply or demand shocks can bring market value and income down as a direct effect, and that households, in response, may be motivated to trim saving in order to maintain consumption. In these cases, the drop in capital growth $g(K)$ would precede and explain the drop in net saving $S_{net}$. The same direction of causality would appear in recoveries from these setbacks. As capital values rise again, that is, income would rise with them, and households would be motivated to restore old saving rates as insurance against the next downturn. (See Appendix E). Net saving, allowing for that margin of possible exception, has raised the physical quantity (volume) of capital, but not the aggregate value, and so has reduced the value per unit. This condition describes capital glut.
@@
The explanation we suggest for capital growth without net saving is in production efficiencies of innovation. The innovator acquires materials and plant capacity and labor skills at market prices determined by their uses in current technology, but applies them more productively until competition catches up. That temporary advantage to the innovator in market value of output per unit cost of input, which we call the innovator’s reserve, creates new income from which to pay that cost, and leaves original income free to pay for consumption. This new income expands the economy, meanwhile, and justifies an equal expansion of the money supply to maintain price stability. Thus growth by a gain in productivity of existing resources costs no sacrifice in consumption, and no diversion of funds from other uses.
An innovator in need of outside funding may approach investors. Investors calculate present value, just as with projects relying on established rather than innovative technology, and invest up to that amount. [^3] Present value is the sum of time-discounted cash flows expected. These flows tend to be negative over any period of investment, and then positive as products are finished and sold. An innovative project may need more investment than others, and a longer period of investment, and will not reward investment unless positive cash flows at the end, discounted to present value, are large enough to offset., [^4] An innovator not in need of outside funding would still require equal present value as a minimum condition for investment., [^5] Competition among investors should tend to keep investment close to present value. Innovative projects may need more investment over a longer period than others, but will still justify funding if expected return at the end is seen as enough to bring present value equal at least to investment. Any excess of present value over investment quantifies the investor’s reserve. [^6] Appendix D will show that an innovative investment will produce cash flows forward at the rate of return $r_{I}$ realized by the innovation, while those cash flows will be discounted backward to present value at the market time-discount rate $r_{m}$ (see Appendix D).
We further suggest, without evidence, that innovation should tend to bring positive externalities to capital of third parties not participating in the innovation. Old structures and infrastructure, for example, should tend to gain value as they house or support higher-tech workers and equipment. We call this effect "tech externalities." "Free growth" will mean the combined growth in capital of innovators through the innovator’s reserve, and in capital generally through tech externalities. Free growth theory will mean the prediction that the rate of productivity gain, which Harrod called the warranted rate, tends to bring an equal rate of capital growth through the innovator’s reserve and tech externalities, without help from net saving, and so brings most or all of capital growth not explained by market noise. Thus the theory allows for some contribution by net saving, but expects any such contribution to be small. We will use the term "thrift residue" to mean any part of capital growth that might be explained by net saving, less than 8% by our findings, with the understanding that the part actually explained by net saving may be zero.
Although net saving is expected to be largely or wholly ineffective, the theory does not question the value of gross saving to make up for losses to depreciation. If these losses are not made up first, value created by productivity gain will go first to make up this shortfall, and only the remainder will apply to capital growth.
@@
The theory predicts only at the largest scales, where the data are clear, and only for the private sector. We accept that households commonly grow or decline in net worth through net saving or dissaving (see Section 16), and that the same is possible for larger groups and even small nations. Free growth theory predicts only at the scale of all capital and all economies together. The public sector, meanwhile, responds to political rather than market choices, and grows or shrinks accordingly.
@@
Inferences for economic teaching include the obvious ones for growth theory and for net saving in general. They include others as well. One of the central doctrines of the marginalist revolution has held that market realization converges to producer cost, when that cost includes imputed interest on assets owned. A gain in productive efficiency can bring market value above producer cost, and by most or all of capital growth if free growth theory holds. Meanwhile the doctrine that net income equals consumption plus net saving is put into question by evidence offered here suggesting that net saving increases the physical quantity of capital, but adds less than 8% to aggregate value.
@@ 2 Net saving, capital growth and capital acceleration
Thrift theory will mean the combined ideas that net saving $S_{net}$ is realized in equal growth of capital $K$, if $S_{net}$ holds within the warranted rate, and that there is no other source of capital growth except market noise regressing to zero. "Thrift assumptions" will mean those two plus the assumption that current $S_{net}$ holds within the warranted rate. The acronym uta will mean "under thrift assumptions." Then
@@
Test results for the predictions $\frac{s^{*}}{g(K)}=1\ \text{and}\ \theta=1$, under thrift assumptions, appear in Fig. 1 and in Tables 1, 2 and 3. Screens to which they refer are explained in Section 15 below. Results were found as GDP-weighted averages over all countries and years, from data for net saving and market-value capital taken from national accounts. Figures 1 and 2 show $\frac{s^{*}}{g(K)}$ and $\frac{\Delta s^{*}}{\Delta g(K)}$ as 0.481 and 0.064 respectively. Table 1 shows regression of $s^{*}$ on $g(K)$ as 0.0771, and regression of $\Delta s^{*}$ on $\Delta g(K)$ as 0.0559.
3 Interpretation of test results
The finding $\frac{s^{*}}{g(K)}=0.481$ reveals average observed proportions between $s^{*}$ and $g(K)$, we repeat, and does not of itself show causal relatedness between those two variables. Causal relatedness is shown rather by the degree of constancy in proportions, and thus on regression of $s^{*}$ on $g(K)$ found at 0.0771. $\theta$ or $\frac{\Delta s^{*}}{\Delta g(K)}$, measured at $0.064$, gives a separate measure of causal relatedness between $s^{*}$ and $g(K)$ as explained. Regression of $\Delta s^{*}$ on $\Delta g(K)$, found at 0.0559, gives a measure of causal relatedness between $\Delta s^{*}$ and $\Delta g(K)$. Might these small but positive findings allow the possibility that some capital growth is explained by net saving?
Appendix E will develop our argument in Section 1 (footnote 2) that the causal relatedness revealed in $\theta$ and its associated regressions might also be explained as joint effects of outside events affecting $g(K)$ directly, and $s^{*}$ indirectly in adjustment to those effects on $g(K)$. We also note that investment in government-owned property such as infrastructure is excluded from predictions of free growth theory, but included in net investment reported in the national accounts from which we take our data. Much government-owned property is unsold, and impractical to appraise, and accordingly is valued in national accounts at depreciated cost. Thus net investment in such government assets shows in national accounts as equal to capital growth, and contributes to $\theta$. Free growth theory takes no position on the actual role of net saving in growth of government-owned property, and consequently neglects this contribution to $\theta$ as outside the scope of its prediction.
4 Inferences for the creative destruction idea
Creative destruction (Schumpeter (1912) [^8] Published by Schumpeter in 1912, and translated by Opie in 1934.) is obsolescence of existing capital through innovation of new and preferred technology, as with horse-drawn carriages and automobiles. Granting the obsolescence of old ways by new, it does not follow that innovation reduces values of assets already present. We have argued that capital growth is maximized when investment is enough to offset depreciation first. A natural source of that investment is depreciation of assets the new technology will replace, and a natural market for the new technology may be owners of the assets to be replaced. Carriage owners were better able to afford cars when the carriages in which they had invested before lived out their economic lives as originally expected, so that the depreciation component in the cash flows originally expected can pay for cars in replacement just as it might have paid for new carriages.
This analysis suggests that market forces favor a kind of orderly obsolescence through replacement, without premature displacement, as the new technology awaits its turn.
Table 1: Regression of $s^{*}$ on $g(K)$ and $\Delta s^{*}$ on $\Delta g(K)$, GDP-weighted. Screen = 0.01. $H_{0}$ per thrift theory: $\frac{s^{*}}{g(K)}\cong\frac{\Delta s^{*}}{\Delta g(K)}\cong 1$.
@@ 4 Inferences for the creative destruction idea
Table 2: Average $\frac{s^{*}}{g(K)}$ in 92 countries over periods shown. Screen = 0.01. Number of years clearing screen shown in ()
@@
Note: $s^{*}$ is defined as $\frac{S_{net}}{K}$. Thrift theory predicts $\frac{s^{*}}{g(K)}\cong 1$. Free growth theory makes no prediction for these data.
Table 3: Average $\theta$ in 92 countries over periods shown. Screen = 0.01. Number of years clearing screen shown in ()
@@
Note: $\theta$ is defined as $\frac{\Delta s^{*}}{\Delta g(K)}$. Thrift theory predicts $\theta\cong 1$. See Section 3 for predictions of free growth theory.
@@ 5 Solow’s puzzle
Solow was right in that much innovation requires months or years of negative cash flow, in the form of gross investment, in building new plant before the new products roll out from the factory doors. From arguments in Section 1, however, this investment would not have been forthcoming unless positive cash flows at the end of the tunnel were predicted to be sufficient to create equal or greater present value from the start.(See Appendix D.) That initial creation of present value, plus replacement investment, [^10] Also called depreciation investment. are enough.
@@ 6 Mechanics of free growth
In terms of Solow’s puzzle, we suggest that embodied growth is disembodied growth on a finer scale. At each step toward realization of the new plant and products, raw materials and products and labor skills and plant capacity currently available on the market are adapted to new uses. The innovator pays for these inputs at a market price determined by their value in established productive uses, but applies them innovatively to realize higher prospective cash flows, and hence higher present values, to the innovator (Marshall (1890), Schumpeter and Opie (1934)). That new use of materials and skills applied could give the solution to Solow’s puzzle with respect to practical feasibility, as argued in Section 1, while the economic advantage to the investor, here called innovator’s reserve,111111i.e., reserve price which the innovator, who realizes more value from capital and labor inputs, might have been willing to pay for those resources. could give the solution with respect to source of payment. Appendix D will clarify.
Our findings support those of Piketty and Zucman (2014) and Kurz (2023) as to the market power of innovators to explain capital growth beyond net saving. Those studies assumed that net saving is recovered in capital growth, so that this market power, meaning the innovator’s reserve, would explain only the part of capital growth remaining. Again, we go farther by questioning the assumption that net saving contributes even a part of capital growth.
@@ 7 Tech Externalities
We argued in Section 1 that innovation can add value eslewhere. Old factories and warehouses and residences gain in utility as they house higher-tech equipment and worker skills. Old highways do the same as they carry higher-tech freight and drivers and passengers to higher-tech destinations in higher-tech vehicles. [^12] Rae (1834) makes essentially the same argument in terms of the technology of his time.
Most externalities, [^13] Externalities are defined as effects of events on nonparticipants. positive and negative, are part of random market noise expected to offset to zero over time and scale. The examples above, which we call tech externalities, are exceptions in that they can impose a prevailingly positive vector on capital value resisting the downward pull of depreciation. Free growth, in the full sense, means effects of the innovator’s reserve on capital of the innovator, plus effects of tech externalities on capital in general.
Note that Solow’s capital deepening, or growth through more efficient use of resources, fits the definition of productivity gain affording the innovator’s reserve. The distinction between capital deepening and widening is otherwise complex. Early investors in technological gain, say venture capitalists, may be rewarded with returns above the current norm. Such early backers form a kind of second-order innovators, and that advantage in return contributes to the innovator’s reserve. Meanwhile other producers may learn and adopt the same technology in a form of externality called spillover, and thus earn above-average returns before the market catches up. Products made by those producers may tend to duplicate those made by the original innovator, but are innovative by comparison to products not yet adopting the innovation. Thus we do not seek to draw a clear line between capital deepening and capital widening, or between tech externalities and the innovator’s reserve.
8 Market value and book value
It is natural to reason that depreciation gives the only change in capital value, once investment has ended, apart from market noise converging to zero. Insofar as depreciation can be predicted, then, depreciated cost would show the norm (expectation) toward which current market value converges. It would follow that investors might anticipate the market by selling when prices are above depreciated cost, and buying when below. This reasoning would justify the reliance of modern national accounts, and the equations of macroeconomics (macro) which they were designed to inform, since the inception of both in Keynes’s time, on depreciated cost as the preferred measure of capital value.  [^14] As argued by James Tobin in 1968 and 1969. ”Tobin’s qq” is the ratio of market to replacement value (RV). RV means depreciated cost under replacement cost accounting, where past investments and subsequent depreciation are re-expressed in the cheaper dollars of today. Investors were advised to invest when Tobin’s $q$ is less than one, and divest or withhold investment conversely.
We have argued that capital grows and declines with effects of the innovator’s reserve and tech externalities, as well as investment and depreciation and market noise. Consider, however, a period when innovation is slight, so that depreciation and market noise give substantially the only changes in value of an investment once booked. It would follow that depreciated cost (of investment) could find the norm described, even if free growth were absent and depreciation were predicted perfectly, only insofar as did original cost. An argument that optimal accounting can give a sounder measure of worth than market value would seem to need a supporting argument that price and worth were closer together at the time assets were acquired or created, and depreciation bases set at their cost, than over the depreciation period following. Garbage in, garbage out.
We do not question the value of depreciation accounting, or the importance of the depreciation concept in economic understanding. But we suggest that current market value gives a more reliable measure of capital, when both are known, than an assumption-dependent projection from past market value subject to the same criticisms as current market value.
These reflections help explain our own reliance on market-value capital. They might also suggest rethinking of the logic by which macro and national accounts have preferred to reason from replacement cost when data for market-value capital are available.